OpenStudy (anonymous):

Find all solutions to the equation. cos^2x + 2 cos x + 1 = 0

OpenStudy (caozeyuan):

you mean\[\cos ^{2}x+2\cos x+1=0 \]

OpenStudy (anonymous):

put Cos x = y and turn the equation into a quadratic in y.

OpenStudy (caozeyuan):

set cos x = A thus A^2+2A=1=0 so A+1)^2 =0 that is A+1=0 meaning A=-1 now substitute back to x giving cos x =-1 can you solve it now? BTW is x is degree or radius and what's the range for value of x?

OpenStudy (anonymous):

it doesnt say if its in radians or degrees @caozeyuan

OpenStudy (caozeyuan):

well I would like to treat it as radius, but you can deal with it whichever way you want

OpenStudy (caozeyuan):

if cos x =-1 that means x=arccos(-1)=pi. if the values are from -pi to pi then, -pi is another solution. if they are supposed to be in 0 to 2pi then there are no other solutions in range.

OpenStudy (dumbcow):

usually default parameters for these problems are : answer in radians from 0 to 2pi